List the three major categories of exceptions and an example of each. Drag the appropriate items to their respective bins 
Answer #1

Octet rule:- According to this rule, central metal atom must contain 8 Electrons. And every atom contribute 8 electrons.

Exception of this rule - (These don't follow octet rule)

• Odd octet

•icomplete octet

• Expanded octet

1) NO:-

Valence electrons of N = 5

Valence electrons of O = 6

Total no.of valence electrons = 5+6 = 11

Electrons contributed by NO through octet rule = 8×2= 16

(Here multiplied with 2 because NO has total No.of atoms =2).

Shared electrons is calculated by-

Shared electrons = Total Electrons contributed through octet rule - total no.of valence electrons

Shared electrons = 16 -11 = 5

As electron sharing can not be odd so odd electrons present on central metal atom.

Hence NO is an example of Odd octet.

2)BCl3 :-

Valence electrons of B = 3

Valence electrons of Cl= 7

Total no. Of valence electrons = 3+(7×3) = 3+21=24

Total valence electrons of BCl3 contributed through octet rule = 8×4=32

(Here multiplied with 4 because BCl3 has total no.of atoms = 4).

Shared electrons = 32 - 24

= 8

But here in Lewis structure it contains only 6 electrons on central metal atom so this is an example of incomplete octet.

Hence BCl3 is an example of incomplete octet.

3) AsF5:

Valence electrons of As =5

Valence electrons of F= 7

Total no.of valence electrons = 5+(7×5) = 5+35=40

Total no of valence electrons contributed through octet rule =

8×6 = 48

(Here multiplied with 6 because AsF5 has total no.of atoms =6).

Shared electrons = 48-40 =8

But here in Lewis structure electrons shared is 10 on central metal atom.

Hence AsF5 is an example of Expanded octet.

:N CL Odd ocie CL comble te ocet ottm has er Be causec

:N CL Odd ocie CL comble te ocet ottm has er Be causec